OpenStudy (anonymous):

Solve by completing the square. x^2 -20x=10

OpenStudy (anonymous):

\[x^2 -20x=10\]

OpenStudy (anonymous):

100

OpenStudy (anonymous):

x^2 -20x +100 = 10 + 100 divide -20 by two and square it.. you add that (100) to both sides.

OpenStudy (anonymous):

knew it had something to do with 100!

OpenStudy (anonymous):

what do I do after that

OpenStudy (anonymous):

(x-10)^2 = 110 \[x-10=\sqrt{110}\] and we get two possible answers for x x=10 + sqrt 110 or x = 10- sqrt 110

OpenStudy (anonymous):

just wanted to say you did an awesome job explaining this. I was able to do the next 10 probs by myself, pretty easily. Thank you :)